OpenStudy (anonymous):

how many committees of four can be chosen from twelve students? how many of these will include a given student? how many will exclude a given student?

OpenStudy (anonymous):

you are being asked for the number of ways to choose 4 from 12, written as \[\dbinom{12}{4}\] read "12 choose 4" and computed via \[\dbinom{12}{4}=\frac{12\times 11\times 10\times 9}{4\times 3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

what about the last two question?

OpenStudy (anonymous):

I don't understand what the last two questions are asking for